# Leveraging Large Language Model for Information Retrieval-based Bug Localization

### v1

@@
Abstract
Information Retrieval-based Bug Localization aims to identify buggy source files for a given bug report. While existing approaches—ranging from vector space models to deep learning models—have shown potential in this domain, their effectiveness is often limited by the vocabulary mismatch between bug reports and source code. To address this issue, we propose a novel Large Language Model (LLM) based bug localization approach, called GenLoc. Given a bug report, GenLoc leverages an LLM equipped with code-exploration functions to iteratively analyze the code base and identify potential buggy files. To gather better context, GenLoc may optionally retrieve semantically relevant files using vector embeddings. GenLoc has been evaluated on over 9,000 real-world bug reports from six large-scale Java projects. Experimental results show that GenLoc outperforms five state-of-the-art bug localization techniques across multiple metrics, achieving an average improvement of more than 60% in Accuracy@1.
Index Terms:
@@ Abstract
I Introduction
@@
Bug localization is the process of identifying the faulty or buggy locations of the source code [1]. It is a crucial step in software debugging to ensure the quality of a software [2]. However, large systems often incur a huge number of bugs everyday. For example, everyday almost 300 bugs are reported in Mozilla, which is overwhelming for the developers [3]. Additionally, given the size and complexity of large-scale software systems today, manually locating bugs is time consuming, tedious and expensive task [4]. As a result, there is a strong demand for automated bug localization techniques to ease the burden of developers and speed up the program debugging process [5].
@@
Over the years, researchers have proposed multiple approaches for bug localization. Among these, Spectrum-based Bug Localization (SBBL) and Information Retrieval-based Bug Localization (IRBL) are among the most studied and widely recognized techniques [6, 5]. SBBL leverages runtime behavior—such as test coverage information—to analyze which parts of a program are exercised by passing and failing test cases, and ranks program elements based on their likelihood of being faulty [7]. In contrast, IRBL takes a bug report—a document that describes a software bug—as input and does not require bug-revealing test cases or execution traces [8]. Due to its simplicity, IR-based approaches have received growing attention in recent years.
@@
IRBL approaches consider the bug report as a query and the software repository as a collection of documents. Given a bug report, these approaches output a ranked list of documents (source files) based on their probability of being faulty. The goal is to rank the actual buggy files as high as possible. For calculating the similarity between bug reports and the source files, earlier techniques use Vector Space Model (VSM) [9, 10], Topic Modeling [11, 12], while more recent methods [13, 14, 15, 16, 17, 18, 5, 19, 20, 21] rely on machine learning and deep learning models. Many of these approaches also incorporate additional features such as historical bug-fix data. However, such information may not always be available, especially in new projects, limiting the generalizability of these methods. Furthermore, the effectiveness of these approaches is often limited by the vocabulary gap between bug reports and source code, as they tend to use different terminologies [6].
Recently, Large Language Models (LLMs) have demonstrated promising results in various software engineering tasks including code generation, code summarization and program repair [22, 23]. Their ability to understand natural language and source code might help improve information retrieval-based bug localization. However, applying LLMs to this task presents unique challenges. Most notably, due to context window limitations, it is infeasible to provide the entire code base as input and ask the LLM to identify relevant files for a given bug report. A straightforward alternative—feeding one file at a time to LLM and asking whether it is relevant—quickly becomes impractical due to the large number of files in real-world repositories. Furthermore, LLMs often struggle to identify and leverage relevant information when the input context is long and relevant information is buried in the middle [24]. Another critical challenge is that LLMs are prone to hallucination (e.g., generating non-existent file names) particularly when they lack sufficient contextual information. Therefore, an effective strategy is needed to help LLMs navigate the code base, selectively retrieve relevant information, and reason within a constrained context window.
@@
In this context, we propose a novel bug localization technique, named GenLoc, that combines semantic retrieval with the step-by-step reasoning capabilities of LLMs to identify buggy files for a given bug report. The name GenLoc stands for Generative AI-based Bug Localization, emphasizing its core design principle of leveraging generative language models for fault localization. Unlike many prior approaches [9, 5], GenLoc relies solely on the bug report and does not require any additional project-specific metadata or historical bug-fix data. At first, GenLoc converts the source code into chunks and measures their semantic similarity with the bug report to retrieve a set of potentially relevant files. After analyzing a bug report, an LLM may choose to inspect these retrieved files further or explore other files if needed. To support this exploration, the LLM is equipped with a set of functions (e.g., retrieving method signatures or method bodies) that enable it to iteratively reason about which files are likely related to the reported bug. This approach allows the LLM to conduct more guided and selective localization.
Figure 1: Workflow of GenLoc
GenLoc is evaluated on over 9,000 real-world bug reports from six large open-source Java projects and compared against five state-of-the-art traditional and deep learning-based IRBL techniques. The results show that GenLoc outperforms all existing approaches across multiple metrics, including Accuracy@1 and Accuracy@5. On average, GenLoc achieves an Accuracy@1 of 43.19%, surpassing all prior techniques by at least 60%. Furthermore, GenLoc offers complementary strengths to existing IRBL techniques, suggesting potential for future integration to further improve bug localization performance.
@@
  • The first study to investigate the impact of incorporating LLMs for information retrieval-based bug localization.
  • The introduction of a novel bug localization approach named GenLoc that integrates semantic retrieval with LLM-guided reasoning.
  • A comprehensive empirical evaluation of GenLoc on over 9,000 real-world bugs from six large open-source Java projects.
  • A publicly accessible replication package [^1] https://github.com/GenLoc-2025/GenLoc to facilitate future research.
II Background
@@
Vector database: A vector database is a special type of database designed to store and manage large collections of high-dimensional embeddings and perform efficient similarity searches [25]. To accelerate the retrieval process, vector databases commonly use Approximate Nearest Neighbor (ANN) algorithms [26], which find vectors that are close to a given query without exhaustively comparing all entries. In GenLoc, the embeddings of all the source files are stored in a vector database to facilitate retrieval of the top source files that are most semantically similar to a given bug report.
Function calling: This mechanism allows LLMs to interact with external tools and systems to retrieve up-to-date information or perform specific tasks that they cannot directly do. For example, if LLM is given the prompt "What is the current temperature in Los Angeles?", it will fail to answer because it lacks access to real-time data. However, it can be provided with access to a function like get_current_temperature(), which queries an external weather API and returns the latest temperature. The model can then call this function, receive the temperature, and integrate it into its response. In GenLoc, the LLM is given access to several functions such as retrieving method signatures and bodies that allow it to iteratively navigate the code base and identify potentially buggy files.
ReAct framework: The ReAct (Reasoning and Acting) framework enables language models to interleave natural language reasoning with tool-based actions, thereby enhancing their problem-solving capabilities [27]. Instead of generating a final answer in a single step, the model alternates between articulating intermediate reasoning steps and performing actions, such as calling a function. This synergy between reasoning and acting allows the model to iteratively refine its understanding, validate hypotheses, and make informed decisions. Since ReAct has shown promise in handling complex tasks that require multi-step thinking, access to external environments, or dynamic decision-making [27, 28], this paper adopts the ReAct framework to guide the LLM in navigating the code base and progressively narrowing down the search space during bug localization.
III Methodology
GenLoc operates in two primary steps to localize relevant files based on a given bug report. First, it retrieves a set of semantically similar files using embedding-based similarity. Next, an LLM, supported by a set of external functions, iteratively analyzes the bug report and source files. It may inspect the embedding-based retrieved files or explore other parts of the code base to generate a ranked list of potentially buggy files. An overview of the workflow is shown in Fig. 1.
III-A Shortlisting suspicious files
@@
The concatenation of a file’s path and method bodies can result in lengthy text. To better localize relevant content, this representation is split into smaller chunks if it exceeds a predefined length threshold [29]. Each chunk is then converted into an embedding vector and stored in a vector database. Similarly, the textual content of the bug report—comprising its summary and description—is transformed into embedding vector.
@@
To measure the similarity between the bug report and the source code, cosine similarity is employed, which measures the angular distance between the embedding vectors [30]. Based on the similarity scores, a predefined number of the top suspicious files are shortlisted. These files serve as candidate inputs for further analysis by the LLM in the subsequent step.
III-B Ranking suspicious files
@@
In this step, an LLM is employed to generate a ranked list of potentially buggy source files based on a given bug report, as illustrated in Fig. 1. To enhance the LLM’s reasoning capabilities, the prompt (shown in Fig. 2) is constructed by following the ReAct (Reasoning and Acting) framework [27], which enables LLMs to interleave natural language reasoning with tool-based actions. This approach has proven effective in improving performance on complex decision-making tasks by enabling iterative reasoning and environment interaction.
At first, the LLM is prompted to assume the role of an expert software engineer specializing in fault localization, as role-playing improves the reasoning abilities of LLMs [31]. Next, the prompt is designed to decompose bug localization into a series of smaller tasks since LLMs perform better when complex tasks are broken down into sub-tasks [32]. These sub-tasks include analyzing the bug report, identifying potential file or method names, analyzing method signatures and bodies, and integrating evidence to prioritize suspicious files.
@@
You are an expert software engineer specializing in fault localization. Your goal is to identify the most probable buggy Java files based on a given bug report. You have access to five functions that will help you infer file names, locate methods, and analyze source code. You must follow an iterative, reasoning-based approach, refining your strategy dynamically based on prior successes and failures. Continue this process until you either (a) produce a well-justified ranked list of the 10 most relevant files based on the bug report, or (b) reach the maximum limit of 10 iterations. In the 10th iteration, you must provide your final output regardless of confidence level. Workflow 1. Analyze the Bug Report: Extract relevant keywords, error messages, and functional hints from the bug summary and description. Identify potential components (e.g., UI, database, networking) involved in the issue. 2. Search: Use search_file() to check if a filename matching the extracted keywords or functionality exists in the code base. If the bug report references a specific method name, use search_method() to locate the file(s) containing that method. If an inferred filename or method location does not exist, refine your strategy: adjust assumptions, explore variations, and retry. If no strong inference can be made, use get_candidate_filenames() to retrieve 50 potential filenames. From the retrieved filenames, prioritize those that align with the bug report’s keywords, functionality, or mentioned methods. 3. Method Analysis: For shortlisted files, retrieve method signatures using get_method_signatures_of_a_file(). Identify methods that directly align with the bug’s context (e.g., matching function names, handling related data). If method signatures suggest a relevant function, retrieve its implementation using get_method_body(). Analyze logic to determine if it aligns with the bug’s symptoms. 4. Refinement and Ranking: Rank files based on multiple factors: Keyword and functionality match Method or filename alignment with bug context Code logic alignment with the bug description If uncertainty remains, refine the analysis by iterating over previous steps with adjusted assumptions. 5. Output: Provide a ranked list of the 10 most relevant files based on their likelihood of containing the bug. Ensure filenames exactly match those provided—do not modify case, structure, or abbreviate them. Justify each file’s inclusion, clearly explaining its relevance to the bug.
Figure 2: LLM Prompt
@@
To support the reasoning process, the LLM is given access to the following five external functions. The LLM may invoke one or more of these functions over a fixed number of iterations to get information about the code base.
  1. 1.
@@
    search_file(): This function enables the LLM to check whether a specific file exists in the code base. It can be used to (i) verify the existence of a file explicitly mentioned in the bug report (e.g., checking if JavaElementLabels.java exists when “JavaElementLabels” is referenced), or (ii) infer and validate a possible filename derived from textual clues in the bug report.
  2. 2.
@@
    search_method(): When a method name is mentioned in the bug report, e.g., updateLabel(), this function enables the LLM to find which files contain its definition and thereby narrowing down the search space.
  3. 3.
@@
    get_candidate_filenames(): This function retrieves a list of potentially relevant files based on their semantic similarity to the bug report from the previous step (Section III-A). The LLM can leverage this list as an initial pool of candidates, especially when the bug report does not provide any clues.
  4. 4.
@@
At each iteration, the model can iteratively “reason” about what information is currently available, determine what additional data is required, and then “act” by invoking one of these functions. The outcome of the function call is integrated into the LLM’s internal reasoning in the next iteration. This iterative loop continues until the model either reaches the maximum allowed number of steps or gathers sufficient information to generate a ranked list of suspicious files.
@@
As LLMs are susceptible to hallucinations, such as generating filenames or method signatures that do not exist in the code base, the system is equipped with recovery strategies to mitigate these issues. For example, when a generated method signature does not match any entry in the code base, approximate candidates are retrieved using fuzzy matching based on the Damerau-Levenshtein distance [33]. Lastly, the LLM is prompted to output a ranked list where each entry includes a fully qualified file path along with a brief justification for its selection. Since multiple files from different packages can have the same name, the fully qualified file path is used to uniquely identify each file. The inclusion of justifications is motivated by prior work demonstrating that self-explanatory prompts significantly improve LLM comprehension [34].
@@
Although several strategies are employed to mitigate LLM hallucinations, such as grounding it through function calls, the model can still generate filenames that do not exist in the code base. To address this, a post-processing step is applied to verify the existence of each predicted file. If the predicted fully qualified file path matches an existing file, it is retained. Otherwise, the base filename (e.g., TitleBlock.java) is extracted, and all files with the same name are retrieved. Among these candidates, the file whose fully qualified file path has the highest Jaccard similarity [35] with the predicted filename is selected. If no suitable match is found, the file is excluded from the final ranked list.
IV Experiement Setup

### v2

@@
Abstract.
Information Retrieval-based Bug Localization (IRBL) aims to identify buggy source files for a given bug report. Traditional and deep-learning-based IRBL techniques often suffer from vocabulary mismatch and dependence on project-specific metadata, while recent Large Language Model (LLM)-based approaches are limited by insufficient contextual information. To address these issues, we propose GenLoc, an LLM-based technique that combines semantic retrieval with code-exploration functions to iteratively analyze the code base and identify potential buggy files. We evaluate GenLoc on two diverse datasets: a benchmark of 9,097 bugs from six large open-source projects and the GHRB (GitHub Recent Bugs) dataset of 131 recent bugs across 16 projects. Results demonstrate that GenLoc substantially outperforms traditional IRBL, deep learning approaches and recent LLM-based methods, while also localizing bugs that other techniques fail to detect.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software maintenance toolsccs: Information systems Information retrieval
1. Introduction
@@
Bug localization is the process of identifying the faulty or buggy locations in source code (Lee et al., 2018). It is a crucial step in software debugging to ensure the quality of a software (Wong and Tse, 2023). However, localizing bugs in practice is challenging, as large systems often incur a huge number of bugs everyday. For example, Mozilla receives around 300 new bug reports daily, which is overwhelming for the developers (Anvik et al., 2006). Moreover, given the size and complexity of large-scale software systems today, manually locating bugs is a time-consuming, tedious and expensive task (Wong et al., 2016). Consequently, there is a strong demand for automated bug localization techniques to ease the burden of developers and accelerate the program debugging process (Qi et al., 2021).
@@
Over the years, researchers have proposed multiple approaches for bug localization. Among these, Spectrum-based Bug Localization (SBBL) and Information Retrieval-based Bug Localization (IRBL) are the most studied and widely recognized techniques (Liu et al., 2024; Qi et al., 2021). SBBL leverages runtime behavior (e.g., test coverage information) to analyze which parts of a program are exercised by passing and failing test cases, and ranks program elements based on their likelihood of being faulty (Zou et al., 2019). In contrast, IRBL takes a bug report—a document that describes a software bug—as input and does not require bug-revealing test cases or execution traces (Kim and Lee, 2021). Due to its simplicity, IR-based approaches have received growing attention in recent years.
@@
IRBL approaches consider the bug report as a query and the software repository as a collection of documents (i.e., source files). Given a bug report, these approaches output a ranked list of source files based on their probability of being faulty. The goal is to rank the actual buggy files as high as possible. To calculate the similarity between bug reports and the source files, earlier techniques use Vector Space Model (VSM) (Zhou et al., 2012; Thomas et al., 2013), Topic Modeling (Lukins et al., 2010; Nguyen et al., 2011), while later methods (Ye et al., 2016; Liu et al., 2019; Lam et al., 2015, 2017; Xiao et al., 2019; Zhang et al., 2020; Qi et al., 2021; Liang et al., 2022; Ye et al., 2015; Chakraborty et al., 2024) rely on machine learning and deep learning models. Many of these approaches also incorporate additional information such as historical bug-fix data. However, such information is not always available—particularly in new projects—thereby limiting the generalizability of these methods. Furthermore, the effectiveness of these approaches is often limited by the vocabulary gap between bug reports and source code, as they tend to use different terminologies (Liu et al., 2024).
To bridge these limitations, recent studies have incorporated Large Language Models (LLMs) into IRBL, as they can reason over both natural language and source code (Samir and Rahman, 2025; Li et al., 2025a). These approaches generally retrieve a small set of candidate files and then use an LLM for further analysis or reranking. For instance, BRaIN shortlists files using lexical similarity (Samir and Rahman, 2025), whereas LLM-BL retrieves candidates using hybrid search (Li et al., 2025a). While such methods improve upon traditional IRBL techniques, their reliance on a restricted context (e.g., a small set of shortlisted files) can result in faulty files being overlooked. If the actual buggy file is absent from the retrieved candidates, the LLM will fail to localize the bug. Expanding the context is a natural alternative, but it introduces new challenges. For example, providing the entire repository as input is infeasible due to context window constraints, and examining files one by one is impractical for large projects. Furthermore, LLMs often struggle to identify relevant information when the input context is long and relevant information is buried in the middle (Liu et al., 2023). These challenges highlight the need for an effective strategy that enables LLMs to navigate the code base, selectively retrieve relevant information and reason within a constrained context window.
@@
In this context, we propose a novel bug localization technique, named GenLoc, that combines semantic retrieval with iterative reasoning capabilities of LLMs to identify buggy files for a given bug report. The name GenLoc stands for Generative AI-based Bug Localization, emphasizing its core design principle of leveraging generative language models for bug localization. Unlike many prior approaches (Zhou et al., 2012; Qi et al., 2021), GenLoc relies solely on the bug report and does not require any additional project-specific metadata or historical bug-fix data. At first, GenLoc converts the source code into chunks and measures their semantic similarity with the bug report to retrieve a set of potentially relevant files. Next, it asks an LLM to analyze a bug report and inspect these retrieved files further, or explore other files as needed. To support this exploration, the LLM is equipped with a set of functions (e.g., retrieving method signatures or method bodies) that enable it to iteratively reason about which files are likely related to the reported bug. This design enables guided and selective localization, avoiding the pitfalls of narrow retrieval that excludes the faulty file and excessive repository-level input that overwhelms the LLM.
We evaluate GenLoc on two complementary datasets: (i) a large benchmark of 9,097 bug reports from six open-source projects that has been widely used in prior bug localization research (Chakraborty et al., 2024), and (ii) the more recent GHRB (GitHub Recent Bugs) dataset of 131 bugs from 16 projects, designed specifically for assessing LLM-based approaches (Lee et al., 2024). Results show that GenLoc consistently outperforms existing approaches. On the large benchmark, GenLoc improves Accuracy@1 by over 63% compared to traditional and deep learning-based IRBL techniques, and on the GHRB dataset, it surpasses recent LLM-based methods by more than 25%. GenLoc also achieves superior ranking performance by placing correct files much earlier in the result list. Moreover, it exhibits complementary strengths to existing techniques, underscoring its potential for integration into hybrid approaches to advance bug localization performance further.
@@
  • The introduction of a novel bug localization approach named GenLoc that integrates semantic retrieval with LLM-guided iterative reasoning.
  • A comprehensive empirical evaluation of GenLoc on more than 9,000 real-world bugs from six large open-source projects and the recent GHRB dataset, including comparisons with traditional IRBL techniques and state-of-the-art LLM-based approaches, as well as an ablation study to assess the contribution of each component.
  • A publicly accessible replication package to facilitate future research.
2. Background
@@
Vector database: A vector database is a special type of database designed to store and manage large collections of high-dimensional embeddings and perform efficient similarity searches (Han et al., 2023b). To accelerate the retrieval process, vector databases commonly use Approximate Nearest Neighbor (ANN) algorithms (Tian et al., 2023), which find vectors that are close to a given query without exhaustively comparing all entries. In GenLoc, the embeddings of all the source files are stored in a vector database to facilitate retrieval of the top source files that are most semantically similar to a given bug report.
Function calling: This mechanism allows LLMs to interact with external tools and systems to retrieve up-to-date information or perform specific tasks that they cannot directly do. For example, if LLM is given the prompt "What is the current temperature in Los Angeles?", it will fail to answer because it lacks access to real-time data. However, it can be provided with access to a function like get_current_temperature(), which queries an external weather API and returns the latest temperature. The model can then call this function, receive the temperature and integrate it into its response. In GenLoc, the LLM is given access to several functions such as retrieving method signatures and bodies that allow it to iteratively navigate the code base and identify potentially buggy files.
ReAct framework: The ReAct (Reasoning and Acting) framework enables language models to interleave natural language reasoning with tool-based actions, thereby enhancing their problem-solving capabilities (Yao et al., 2023). Instead of generating a final answer in a single step, the model alternates between articulating intermediate reasoning steps and performing actions, such as calling a function. This synergy between reasoning and acting allows the model to iteratively refine its understanding, validate hypotheses, and make informed decisions. Since ReAct has shown promise in handling complex tasks that require multi-step thinking, access to external environments, or dynamic decision-making (Yao et al., 2023; Wu, 2025), this paper adopts the ReAct framework to guide the LLM in navigating the code base and progressively narrowing down the search space during bug localization.
Figure 1. Workflow of GenLoc.
3. Methodology
GenLoc operates in two primary steps to localize relevant files based on a given bug report. First, it retrieves a set of semantically similar files using embedding-based similarity. Next, an LLM, supported by a set of external functions, iteratively analyzes the bug report and the code base. During this stage, the model may examine the embedding-based retrieved files or explore other parts of the code base to generate a ranked list of potentially buggy files. An overview of the workflow is shown in Fig. 1.
3.1. Shortlisting suspicious files
@@
The concatenation of a file’s path and method bodies can result in lengthy text. To better localize relevant content, this representation is split into smaller chunks if it exceeds a predefined length threshold (Wong et al., 2014). Each chunk is then converted into an embedding vector and stored in a vector database. Similarly, the textual content of the bug report—comprising its summary and description—is transformed into embedding vector.
@@
To measure the similarity between the bug report and the source code, cosine similarity is used, which measures the angular distance between the embedding vectors (Steck et al., 2024). Based on the similarity scores, a predefined number of the top suspicious files are shortlisted. These files serve as candidate inputs for further analysis by the LLM in the subsequent step.
3.2. Ranking suspicious files
@@
In this step, an LLM is employed to generate a ranked list of potentially buggy source files based on a given bug report, as illustrated in Fig. 1. To enhance the LLM’s reasoning capabilities, a prompt (shown in Fig. 2) is constructed by following the ReAct (Reasoning and Acting) framework (Yao et al., 2023), which enables LLMs to interleave natural language reasoning with tool-based actions. This approach has proven effective in improving performance on complex decision-making tasks by enabling iterative reasoning and environment interaction.
@@
You are an expert software engineer specializing in bug localization. Your goal is to identify the most probable buggy Java files based on a given bug report. You have access to five functions that help you search for file names, locate methods and analyze source code. You must follow an iterative, reasoning-based approach, refining your strategy dynamically based on insights gathered during each step. At each iteration, you should: Maintain a working shortlist of files that appear potentially buggy. Update the shortlist based on new evidence (e.g., method matches, code analysis). Avoid redundant operations–do not recheck the same filenames, methods or file contents multiple times. Continue this process until you either: (a) Produce a well-justified ranked list of the 10 most relevant files, or (b) Reach the maximum limit of 10 iterations. In the 10th iteration, you must output your final ranked list. Workflow 1. Analyze the Bug Report: Extract relevant keywords, error messages and functional hints from the bug summary and description. Identify potentially affected components (e.g., UI, database, networking). 2. File Discovery: Use search_file() to check if filenames derived from the bug report’s keywords or functionality exist in the code base. If the bug report references a specific method name, use search_method() to find all files defining it. If an inferred filename or method location does not exist, refine your strategy: adjust assumptions, explore variations and retry. If strong matches are not found, use get_candidate_filenames() to retrieve 50 potentially relevant filepaths. Add promising files to your shortlist by prioritizing those align with terminology, functionality or methods discussed in the bug report. 3. Method Analysis: For each file in the shortlist that hasn’t been analyzed yet: Use get_method_signatures_of_a_file() to list its methods. Identify methods that directly align with the bug’s context (e.g., related functionality, naming hints). For any method of interest, retrieve its implementation using get_method_body(). Analyze the logic of any identified method(s) of interest to determine whether they align with the bug’s symptoms. If this analysis reveals new relevant class names, filenames, or methods, use the appropriate search functions. Continuously update the shortlist by promoting, demoting or removing files based on evolving understanding. 4. Shortlist Refinement & Ranking: Rank files based on: Semantic alignment with the bug report’s keywords and described functionality Method or filename alignment with bug context Code logic alignment with the bug description If needed, iterate with refined assumptions or explore previously overlooked filenames or methods. 5. Final Output: Provide a ranked list of the 10 most relevant filepaths based on their likelihood of containing the bug. Ensure filepaths exactly match those provided–do not modify case, structure or abbreviate them. Justify each file’s inclusion by referencing keywords, method matches or code logic that supports its relevance to the bug report.
Figure 2. LLM Prompt.
At first, the LLM is prompted to assume the role of an expert software engineer specializing in bug localization, as role-playing improves the reasoning abilities of LLMs (Kong et al., 2023). Next, the prompt is designed to decompose bug localization into a series of smaller tasks since LLMs perform better when complex tasks are broken down into sub-tasks (Khot et al., 2022). These sub-tasks include analyzing the bug report, identifying potential file or method names, analyzing method signatures and bodies and integrating evidence to prioritize suspicious files.
@@
To support the reasoning process, the LLM is given access to the following five external functions. The LLM may invoke one or more of these functions over a fixed number of iterations to get information about the code base. Collectively, these functions allow the LLM to discover files, understand their role and analyze the internal logic of specific code segments, thereby supporting both breadth and depth in the bug localization process.
  1. (1)
@@
    search_file(): This function enables the LLM to check whether a specific file exists in the code base. It can be used to (i) verify the existence of a file explicitly mentioned in the bug report (e.g., checking if JavaElementLabels.java exists when “JavaElementLabels” is referenced), (ii) infer and validate a possible filename derived from textual clues in the bug report, or (iii) identify candidate files after inspecting a method body.
  2. (2)
@@
    search_method(): When a method name is mentioned in the bug report, e.g., updateLabel(), this function enables the LLM to find which files contain its definition and thereby narrowing down the search space. In addition, it can be employed to discover candidate files based on references observed during method body inspection.
  3. (3)
@@
    get_candidate_filenames(): This function retrieves a list of potentially relevant files based on their semantic similarity to the bug report from the previous step (Section 3.1). The LLM can leverage this list as an initial pool of candidates, especially when the bug report does not provide any clues.
  4. (4)
@@
At each iteration, the model can “reason” about what information is currently available, determine what additional data is required, and then “act” by invoking one of these functions. The outcome of the function call is integrated into the LLM’s internal reasoning in the subsequent iteration. Since the LLM can make mistakes while invoking function calls, such as supplying incorrect parameters like non-existent filenames or method signatures, we incorporate an iteration-time recovery step. When no exact match is found, the system either provides tentative options (e.g., all file paths sharing the same base filename or the closest method signatures using Damerau–Levenshtein distance (Zhao and Sahni, 2019)) or returns explicit feedback indicating that the requested element does not exist in the code base. This mechanism guides the LLM in handling mistakes (e.g, typos or incomplete names) by recovering the intended file or method, or preventing it from producing non-existent ones.
@@
Lastly, the LLM is prompted to output a ranked list where each entry includes a fully qualified file path along with a brief justification for its selection. Since multiple files from different packages can have the same name, the fully qualified file path is used to uniquely identify each file. The inclusion of justifications is motivated by prior work demonstrating that self-explanatory prompts significantly improve LLM comprehension (Wu et al., 2023).
@@
Although the LLM receives feedback during intermediate iterations, it can still produce erroneous or non-existent file paths in the final ranked list. A common issue we observed in our preliminary analysis is the omission of intermediate packages. For example, the model may omit the "internal" package and produce org.eclipse.jdt.ui/ui/org/eclipse/jdt/ui/text/java/AlphabeticSorter.java instead of org.eclipse.jdt.ui/ui/org/eclipse/jdt/internal/ui/text/java/AlphabeticSorter.java. To address this, we apply a post-processing recovery step to verify the correctness of each predicted file. If the predicted fully qualified file path matches an existing file, it is retained. Otherwise, the base filename (e.g., AlphabeticSorter.java) is extracted, and all files with the same name are retrieved. Among these candidates, the file whose fully qualified file path has the highest Jaccard similarity (Niwattanakul et al., 2013) with the predicted filename is selected. Since Jaccard measures token overlap, it favors the correct file even when an intermediate package is missing. If the base filename does not exist in the code base, no suitable match can be found and the file is excluded from the final ranked list.
4. Experiement Setup